Last month, the town of New Hen established a parking enforcement unit (PEU) to write parking citations. New Hen budgeted $3600 for the unit's first 30 days, and estimated PEU officers would write 50 citations per day in those 30 days. In addition, New Hen estimated that officers would work 12 minutes per citation written. New Hen also paid PEU officer $.40 more per hour than budgeted. Calculate the volume, quantity, price, and total expense variances, and indicate whether they are favorable or unfavorable.
Answer #1

Total citation required in 30 days= 50*30=1500

budgeted cost per citation=3600/1500= $2.4

Total working hour=12*1500/60=300 hour

Budgeted cost per hour=3600/300=$12/hour

actual cost per hour=(12+0.4)=$12.4/hour

Total actual cost =12.4*300=$3720

Hence, total volume=1500 citation

price=$3720

expense variance=$12.4 per hour for 300 hour

As the actual cost is more than budgeted cost, this is unfavorable.